A car is emmiting sound of frequency 520 Hz. You are on a motorcylce, traveling directly away from the parked car.
Acaram is om ting sound waves of treguency 520 Ha You are on a motorcycle traveling directly from the parked car Part A What what is your speed if you detect a frequency of 490 Hz?
Answer #1

speed of bike, f = f (V-vo) ,, w - 490= 520(343-vo) 343 Vo= 19.788 mis iere, f=52042 fl=49042 Ov=343 mis